# Biomedical open source software: Crucial packages and hidden heroes

### v3

@@ 1 Introduction
Fig 1: XKCD on software dependency and ecosystem fragility [1]. Also has roll-over text of Someday ImageMagick will finally break for good and we’ll have a long period of scrambling as we try to reassemble civilization from the rubble. Used with permission: https://creativecommons.org/licenses/by-nc/2.5/
@@
Besides this, there is another kind of invisibility of scientific software. The programs visible to the end user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below. These may be packages that the utilized program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. This complex nature of the network of dependencies has a number of implications, including for security [13] and computational reproducibility [14]. In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
The situation resembles the famous XKCD cartoon shown in  Figure 1, where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1]. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termedtransitive credit [15, 16] but so far the problem has not yet been solved. To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [17]. Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [18].
@@ 3 Results
Table 1: Summary statistics for the number of mentions in papers. “Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Gini” is Gini coefficient for the mentions count.
@@
Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, “Gini” is Gini coefficient.
@@
Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLoS CompBio for suggesting this metric). It is shown in Table 1.
In Tables 3, 4, and 5 we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
@@ 4 Discussion
The difference between the unweighted and weighted versions of our analysis, in Figure 4 can help us parse different subtypes of these outliers. For example, the velvet [40] package is central solely because it is mentioned by one highly cited paper (and dozens of less highly cited ones). This strategy is not central in the unweighted version of the analysis, since citation counts are ignored. Central packages under this representation, therefore, also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
@@
    Our approach is limited to the open-source packages available in one of the chosen ecosystems. This excludes proprietary software like Excel or GraphPad Prism, and open-source non-package software like Gephi [43] (a complied application). The number of missing packages may be estimated from Table 14 of [11], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [44]. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services). Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers111as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [45].

### v4

@@ 1 Introduction
Besides this, there is another kind of invisibility of scientific software. The programs visible to the end user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below. These may be packages that the utilized program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. This complex nature of the network of dependencies has a number of implications, including for security [Hatta2022Nebraska, 13] and computational reproducibility [14]. In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
The situation resembles the famous XKCD cartoon where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1]. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termed transitive credit [15, 16] but so far, the problem has not yet been solved. To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [17]. Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [18].
@@ 3 Results
Table 1: Summary statistics for the number of mentions in papers. “Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Max” is the maximal number of mentions, and “Gini” is the Gini coefficient for the mentions count.
@@
Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, and “Gini” is the Gini coefficient.
@@
Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLOS Computational Biology for suggesting this metric). It is shown in Table 1.
In Tables 3, 4, and 5, we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
@@ 4 Discussion
The difference between the unweighted and weighted versions of our analysis, in Figure 3 can help us parse different subtypes of these outliers. For example, the velvet [40] package is central solely because it is mentioned by one highly cited paper (and dozens of less highly cited ones). Such characteristics do not make a package central in the unweighted version of the analysis, since citation counts are ignored there. Central packages under this representation, therefore, also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
@@
    Our approach is limited to the open-source packages available in one of the chosen ecosystems. This excludes proprietary software like Excel or GraphPad Prism, and open-source non-package software like Gephi [43] (a complied application). The number of missing packages may be estimated from Table 14 of [11], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [44]. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools [Hatta2022Nebraska]. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services). Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers111as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [45].